# Performance of Zero-Shot Time Series Foundation Models on Cloud Data

### v2

@@ 3 Experiments
Experimental setup   We evaluate a suite of FMs on cloud time series data. We examine several of the most recent and well know FMs for time series: VisionTS (Chen et al., 2024), TTM (revision 2) (Ekambaram et al., 2024), TimesFM (Das et al., 2024), Chronos (tiny) (Ansari et al., 2024), Moirai (small) (Woo et al., 2024) and Mamba4Cast (Bhethanabhotla et al., 2024). We look at their performance on datasets constructed out of data released on Huawei Cloud function requests, [^1] Available at https://github.com/sir-lab/time-series-fm-dataset as described in Appendix A.1. We name the datasets D1, D2, D3 and D4, each of which correspond to data from a unique data centre in the Huawei Cloud. We note that, none of the FMs looked at have been pretrained on this data or cloud demand data more generally. As explained in Section 2, we perform experiments using a rolling window evaluation. This is done with a context length of $520$ and for three different forecast horizons: $30,96,336$. We selected these values as they have been widely used in previous works (Ekambaram et al., 2024; Ansari et al., 2024).
Baselines   We compare the FMs against two simple baselines: a) A per-channel online linear model which is fit using ridge regression (Hamilton, 1994). This model is refit on all available data every $200$ time steps. b) A naive seasonal forecaster which forecasts by copying repeatedly the last seasonal period in the context as its forecast (Ansari et al., 2024). For example, for hourly data with a daily periodicity, the naive seasonal forecaster would predict by taking the last 24 hours and repeating it over the forecast window. We note that the naive seasonal forecaster does not need to have access to any training data. More details on the baselines can be found in Appendix A.2.
Metrics   We report the MASE of each approach, a common time series evaluation metric (Ansari et al., 2024). MASE is the mean absolute error (MAE) normalized by dividing by the MAE of a naive seasonal forecaster over the context, as detailed in Appendix A.4. This normalisation adjusts for varying scales across channels and over time. We also report results using RMSSE in Appendix B.4, which is defined analogously to MASE but with RMSE instead of MAE (Hyndman & Koehler, 2006). Both metrics yield consistent conclusions.
@@
Results   The main results of our experiments are displayed in Table 1. The table shows that none of the FMs being evaluated perform as well as either of the simple baselines we compare to. For example, the naive seasonal forecaster performs better than all the FMs across all datasets and forecast horizons. Moreover, the performance difference is often large; for example, the naive seasonal forecaster incurs a MASE typically half that of TimesFM. This suggests that currently zero-shot FMs do not generalise well to cloud data. We also note that alongside the good predictive performance of the baseline methods they are also the most computationally efficient, as demonstrated in Appendix B.1.
@@ 3.1 What do Zero-Shot FM Forecasts Look Like?
To analyse further the poor performance of zero-shot FMs in the cloud domain we plot a typical failure case of FMs in Figure 2. The figure displays the forecasts for TimesFM, TTM, Mamaba4Cast and Chronos on the third channel of the D2 dataset—forecasts for Moirai and VisionTS are given in Appendix B.2. The context (shown by the blue curve) is standard of cloud time series, having a strong seasonal pattern and large spikes (Diao et al., 2024), and the ground truth is roughly a repeat of the seasonal pattern. However, despite the regularity of the time series, all of the FMs give poor predictions, appearing to not make use of the clearly periodic behaviour. For example, three of the FMs, in Figure 2, do not identify the seasonal pattern at all; while the other, TTM, gets the rough shape correct but misses the spikes. We note that, in cloud provision settings predicting demand spikes is vital, a failure to do so can degrade service quality.
@@
Overall, the example forecasts in Figures 1 and 2 show that for typical cloud data, zero-shot FM forecasts can fail to resemble the ground truth, even when the ground truth is readily predictable from the context given—as demonstrated by the baseline forecasts given in Appendix B.2. We emphasise that the failures shown are not unique to the displayed channels or time steps, being commonplace in our experiments, evidencing the poor overall performance of FMs on cloud data shown in Table 1.
VisionTS   often performs quite well relative to the other FMs in our experiments (see Table 1). While its performance is still worse than the baselines, it is worth exploring why this is. We find in our explorations that VisionTS behaves very similarly to a naive seasonal forecaster in most cases, simply coping the context as its forecast. An example of this behaviour is shown in Figure 3. Since most channels are relatively periodic this strategy performs well, largely explaining the good performance of this model. To analyse this further, we computed the normalised (zero-lag) cross-correlation between the forecasts of VisionTS and the naive seasonal forecaster on D1 obtaining a median correlation of 0.9992.
@@ 4 Conclusions
Appendix A Additional Experimental Details
A.1 Construction of the Huawei Cloud Datasets
@@ A.4 Metrics
Appendix B Additional Experiments
B.1 Computational Cost
@@
B.2 Baseline and Other FM Forecasts for Figure 2

### v3

@@ 3 Experiments
Experimental setup   We evaluate a suite of FMs on cloud time series data. We examine several of the most recent and well know FMs for time series: VisionTS (Chen et al., 2024), TTM (revision 2) (Ekambaram et al., 2024), TimesFM (Das et al., 2024), Chronos (tiny) (Ansari et al., 2024), Moirai (small) (Woo et al., 2024) and Mamba4Cast (Bhethanabhotla et al., 2024). We look at their performance on datasets constructed out of data released on Huawei Cloud function requests, [^1] Available at https://github.com/sir-lab/time-series-fm-dataset as described in Appendix B.1. We name the datasets D1, D2, D3 and D4, each of which correspond to data from a unique data centre in the Huawei Cloud. We note that, none of the FMs looked at have been pretrained on this data or cloud demand data more generally. As explained in Section 2, we perform experiments using a rolling window evaluation. This is done with a context length of $520$ and for three different forecast horizons: $30,96,336$. We selected these values as they have been widely used in previous works (Ekambaram et al., 2024; Ansari et al., 2024).
Baselines   We compare the FMs against two simple baselines: a) A per-channel online linear model which is fit using ridge regression (Hamilton, 1994). This model is refit on all available data every $200$ time steps. b) A naive seasonal forecaster which forecasts by copying repeatedly the last seasonal period in the context as its forecast (Ansari et al., 2024). For example, for hourly data with a daily periodicity, the naive seasonal forecaster would predict by taking the last 24 hours and repeating it over the forecast window. We note that the naive seasonal forecaster does not need to have access to any training data. More details on the baselines can be found in Appendix B.2.
Metrics   We report the MASE of each approach, a common time series evaluation metric (Ansari et al., 2024). MASE is the mean absolute error (MAE) normalized by dividing by the MAE of a naive seasonal forecaster over the context, as detailed in Appendix B.4. This normalisation adjusts for varying scales across channels and over time. We also report results using RMSSE in Appendix C.4, which is defined analogously to MASE but with RMSE instead of MAE (Hyndman & Koehler, 2006). Both metrics yield consistent conclusions.
@@
Results   The main results of our experiments are displayed in Table 1. The table shows that none of the FMs being evaluated perform as well as either of the simple baselines we compare to. For example, the naive seasonal forecaster performs better than all the FMs across all datasets and forecast horizons. Moreover, the performance difference is often large; for example, the naive seasonal forecaster incurs a MASE typically half that of TimesFM. This suggests that currently zero-shot FMs do not generalise well to cloud data. We also note that alongside the good predictive performance of the baseline methods they are also the most computationally efficient, as demonstrated in Appendix C.1.
@@ 3.1 What do Zero-Shot FM Forecasts Look Like?
To analyse further the poor performance of zero-shot FMs in the cloud domain we plot a typical failure case of FMs in Figure 2. The figure displays the forecasts for TimesFM, TTM, Mamaba4Cast and Chronos on the third channel of the D2 dataset—forecasts for Moirai and VisionTS are given in Appendix C.2. The context (shown by the blue curve) is standard of cloud time series, having a strong seasonal pattern and large spikes (Diao et al., 2024), and the ground truth is roughly a repeat of the seasonal pattern. However, despite the regularity of the time series, all of the FMs give poor predictions, appearing to not make use of the clearly periodic behaviour. For example, three of the FMs, in Figure 2, do not identify the seasonal pattern at all; while the other, TTM, gets the rough shape correct but misses the spikes. We note that, in cloud provision settings predicting demand spikes is vital, a failure to do so can degrade service quality.
Overall, the example forecasts in Figures 1 and 2 show that for typical cloud data, zero-shot FM forecasts can fail to resemble the ground truth, even when the ground truth is readily predictable from the context given—as demonstrated by the baseline forecasts given in Appendix C.2. We emphasise that the failures shown are not unique to the displayed channels or time steps, being commonplace in our experiments, evidencing the poor overall performance of FMs on cloud data shown in Table 1.
@@
VisionTS   often performs quite well relative to the other FMs in our experiments (see Table 1). While its performance is still worse than the baselines, it is worth exploring why this is. We find in our explorations that VisionTS behaves very similarly to a naive seasonal forecaster in most cases, simply coping the context as its forecast. An example of this behaviour is shown in Figure 3. Since most channels are relatively periodic this strategy performs well, largely explaining the good performance of this model. To analyse this further, we computed the normalised (zero-lag) cross-correlation between the forecasts of VisionTS and the naive seasonal forecaster on D1 obtaining a median
correlation of 0.9992
@@ 4 Conclusions
Acknowledgments
We would like to kindly thank Luke Darlow for his help and Thomas L. Lee would like to thank his PhD supervisor, Amos Storkey, for his general guidance on research practice.
Appendix A Generalization in Zero-Shot Time Series Forecasting
While in the main text we have aimed to show if and how zero-shot FMs fail to forecast cloud data here we examine the question of why. To do this we ask more generally how much can we hope to generalize in time series forecasting? In both the vision and text domain it has been shown that zero-shot FMs have good generalization capabilities (Minaee et al., 2024; Zhou et al., 2022) and hence it can be natural to suggest by using zero-shot FMs in the time domain we can achieve similar generalization capabilities. However, there are reasons to question this belief, mainly due to less constrained nature of the continuation of a time series versus the continuation of an image (i.e. inpainting) or text generation. For instance, given the same context there maybe many valid forecasts based on the time series domain the data was generated from. More concretely, in cloud data there are spikes which happen at regular intervals which should be forecasted but in other domains spikes could be caused by noise and so should not appear in the forecast. This suggests that as the claims of zero-shot time series FM generalization expands it is quickly apparent that there is no free lunch. Hence, claims of widespread generalization ability or generalization to untested domains for zero-shot forecasts should be inspected carefully.
Additionally, the above suggest that to create well performing FMs for time series data there needs to be a mechanism for conditioning on previous data from the time series being forecasted or from similar time series. This can be straightforwardly achieved by finetuning methods. However, this can be compute and time intensive and requires access to a significant amount of offline data. So, in real-world scenarios where time series data is generated online, shifts in distribution and compute/training time is constrained, there is a need to look at new conditioning mechanisms. This has already been looked at in some works (Lee et al., 2025; Ma & Fildes, 2023) but there is scope to do more in this area which could realize the goal of general time series FMs.
Appendix B Additional Experimental Details
B.1 Construction of the Huawei Cloud Datasets
@@ B.4 Metrics
Appendix C Additional Experiments
C.1 Computational Cost
@@
C.2 Baseline and Other FM Forecasts for Figure 2